a) Differentiate between purchased goodwill and internally generated goodwill. (5 marks)


b) Upendo ltd wishes to take advantage of the new commercial paper now popular. It


wishes to issue two debenture papers. Both bear coupons of 12% and the effective yield


required on each is 20%. Paper A has a maturity of 10 years and paper B a maturity of


20 years. Both will be paying interest annually and Kshs. 100,000 at maturity. What is


the price of each paper? If the effective yield on each paper rises to 24%, what is the


price of each paper? (15 mks)

1
Expert's answer
2021-11-02T16:59:34-0400

a)

Purchased Goodwill arises from the acquisition of an existing business, while non-purchased goodwill has been built-up over time and cannot be verified objectively.


b)

Price = Face Value/ [1 + yield x (no. of days to maturity/365)]


Face Value="100000\/(1+0.12)^n"


for paper A:

Face Value="100000\/(1+0.12)^{10}=32197.32"

Price = "32197.32/(1+0.2\cdot10)=10732.44"


for paper B:

Face Value="100000\/(1+0.12)^{20}=10366.68"

Price = "10366.68/(1+0.2\cdot20)=2073.34"


If the effective yield on each paper rises to 24%:


for paper A:

Price = "32197.32/(1+0.24\cdot10)=9469.80"


for paper B:

Price = "10366.68/(1+0.24\cdot20)=1787.36"